OpenStudy (anonymous):

The manager of a department store wants to build a 600 square-foot rect- angular enclosure in the stores parking lot to display some equipment. Three sides of the enclosure will be built of redwood fencing at a cost of $14 per running foot. The forth side will be built with cement blocks, at a cost of $28 per running foot. Find the dimensions of the enclosure that will minimize the total cost of the building materials. [Hint: You’ll need to find an equation for the total cost of the materials in terms of only one variable.]

hero (hero):

1. The total area in sq. ft. of the store parking lot in terms of dimensions x and y is: xy = 600 2. The perimeter of the the lot in terms of x and y is: P(x,y) = 3x + y 3. The total cost of the lot in terms of x and y is: C(x,y) = 42x + 28y 4. The total cost of the lot in terms of x only is: C(x) = 42x + 28(600/x) C(x) = 42x + 16800/x C(x) = 42x^2/x + 16800/x C(x) = (42x^2 + 16800)/x 5. The derivative of C'(x) is: C'(x) = 42(x^2 - 400)/x^2 6. Set C'(x) = 0 0 = 42(x^2 - 400)/x^2 0 = 42(x^2 - 400) 0 = x^2 - 400 400 = x^2 x = 20 y = 30 7. Thus the dimensions, (x,y), of the enclosure that will minimize the total cost of building materials is (20,30) in feet.